OpenStudy (anonymous):

Solve each formula for the indicated letter. Assume that all variables represent nonnegative numbers. w=√1/LC , for L

OpenStudy (anonymous):

Then underneath it says (an electricity formula)

OpenStudy (anonymous):

w=sqrt(1/LC) square both sides: w^2=1/LC multiply both sides by LC: LCW^2=1 divide both sides by CW^2: L=1/(CW^2)